OpenStudy (anonymous):

The length of a rectangle is 5ft less than three times the width, and the area of the rectangle is 50ft^2. Find the dimensions of the rectangle. L*W L*W=50 L=3w-5 (3w-5)*W=50 L=5?

OpenStudy (anonymous):

@agent0smith

OpenStudy (agent0smith):

L = 3W - 5 Area = L*W = W(3W-5) = 50 W(3W-5) = 50 distribute and solve for W, prob using quadratic formula

OpenStudy (anonymous):

I meant w=5

OpenStudy (anonymous):

and L = 10

OpenStudy (agent0smith):

Yep :)